Leo who has a mass of 70 kg stretches

the trampoline mat by 37 cm. Calculate the

elastic potential energy he puts into the

trampoline mat.


1
Expert's answer
2022-03-15T10:46:23-0400

Answer

Spring constant

"K=\frac{mg}{x}\\=\frac{70*9.81}{0.37}\\=1.85*10^3N/m"

So

elastic potential energy

"U=\frac{kx^2}{2}\\=\frac{1.85*10^30.37*0.37}{2}\\=127.04J"